Solve by Cramer's Rule x=dx/d, y=dy/d, and z=dz/d
2x-y+3z=9
x+4y+4z=5
3x+2y+2z=5
I have worked this problem and I know it is not right because of the numbers I am getting. I do have that D is -40. I have dx is -20. I have dy is 210 and dz is 156. Something is not right in the way I set up teh matrix. Thanks for your help.

Answer by Alan3354(69443)

D is -40
----------
For Dx:
|9 -1 3|
|5 +4 4|
|5 +2 2|
= 9(8-8) + 1(10-20) + 3(10-20)
= -10 -30
= -40
x = +1
------
For Dy:
|2 9 3|
|1 5 4|
|3 5 2|
= 2(10-20) - 9(2-12) + 3(5-15)
= -20 +90 -30 = +40
y = -1
-------
For Dz:
|2 -1 9|
|1 +4 5|
|3 2 5|
= 2(20-10) + 1(5-15) + 9(2-12)
= 20 - 10 -90
= -80
z = 2